The dimension of a rectangle are such that it's length is 5 in. more than its width. If the length were doubled and if the width were decreased by 5 in. the area would be increased by 162 in^2. What are the length and width of the rectangle?
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

width=w
length=w+5 units inches
Area, A1= w(w+5)=w^2+5w
A2=2(w+5)(w-5)=2w^2-50
w^2+5w+162=2w^2-50
0=w^2-5w-212
w=(1/2)(5+/- sqrt (25+848); sqrt 873=29.55
positive root is w=17.28 inches ANSWER
length is 22.28 inches ANSWER
area is 385
double length and get 44.56
decrease width by 5 and get 12.28
area is 547.20, 162 more with rounding error.

Answer by Alan3354(69443)

it's = it is